OpenStudy (anonymous):

factoring 110 y^2+17y+6

OpenStudy (anonymous):

discriminant is negative

OpenStudy (anonymous):

didnt get

OpenStudy (anonymous):

please sovle

OpenStudy (anonymous):

when u have a quadratic and u want to factor it...actually u going to rearrange it in the form\[c(y-a)(y-b)\]where \(a\) and \(b\) are roots of quadratic...but here discriminant is negative so u have 2 complex roots...

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

when u have a quadratic like this :\[x^2-5x+6\]factoring is easy because roots are integer numbers ... so easily u say\[-2-3=-5\]\[(-2)(-3)=6\]so u write\[x^2-5x+6=(x-2)(x-3)\]